# Fast Global Localization on Neural Radiance Field

### v1

@@ 4.2 Localization on Multi-Scale Representation
Table 2: Comparison on LLFF dataset [36] using Zip-NeRF [13] as a map.
@@
Experiment Setup: We utilized multi-scale representation via Zip-NeRF [13] to conduct comparative analyses between Loc-NeRF and Fast Loc-NeRF. The experimental dataset consists of all 8 scenes from the LLFF dataset [36], consistent with Section 4.1, and is extended to include two additional scenes from the Deep Blending dataset [37]: $playroom$ and $Drjohnson$. The Deep Blending dataset [37], unlike the LLFF dataset [36] which comprises primarily forward-facing scenes, showcases entire rooms, making it particularly suitable for demonstrating global localization for the large real world.
Evaluation on LLFF dataset We present the results of employing Zip-NeRF [13] as a map on our method and Loc-NeRF in Table 2. Our approach, Fast Loc-NeRF demonstrated superior performance against baseline across most of the forward-facing scenes. While Loc-NeRF [10] had difficulty reaching a 0.8 accuracy threshold in any scene, our Fast Loc-NeRF method consistently attained this benchmark across all scenes except for $fern$ and $orchids$ scenes, still demonstrating the highest performance.
@@
Evaluation on Deep Blending dataset [37] In the evaluation conducted on the Deep Blending dataset, we observed that each scene, representing a single room, presents a more challenging environment for global localization compared to the forward-facing scenes in the LLFF dataset. Our method significantly outperforms Loc-NeRF shown on Table 3. This enhancement suggests that our method efficiently understands and explores extensive spatial environments, such as larger rooms, indicating its adaptability and efficiency in addressing more complex localization challenges.
Drjohnson Playroom Average
Loc-NeRF 0.83/54.40 0.97/35.65 0.90/45.03
Ours 0.42/34.04 0.37/33.40 0.40/33.72
(a) Position (cm)/Rotation (deg) Error
Drjohnson Playroom Average
Loc-NeRF 0.10/0.10 0.10/0.15 0.10/0.13
Ours 0.35/0.25 0.25/0.40 0.30/0.33
(b) Position/Rotation Accuracy
Table 3: Comparison on Deep Blending Dataset [37] using Zip-NeRF [13] as a map.
5 Conclusion
In summary, we introduce Fast Loc-NeRF, which exploits particle filtering in a coarse-to-fine manner to enable more efficient and accurate localization on a NeRF map. Leveraging coarse-to-fine matching, Fast Loc-NeRF speeds up the expansive particle weight update process, while progressively matching particle states and measurements from low to high resolution. Furthermore, to deal with particles distributed in irrelevant locations, we propose particle rejection weighting, which estimates ray uncertainty by analyzing the distribution over each ray without any additional training step. By not simply using the RGB values obtained from rendering to weight the particles required for the update step, but also applying the uncertainty of each ray, we reduced the exploration of inappropriate space, enabling faster localization. As a result, our method speeds up the localization process by 55% and improves localization accuracy by 19% with the evaluation of various NeRF datasets and multi-scale NeRF models.
Limitation and Future Work Our Fast Loc-NeRF improves both the performance and efficiency of global localization on NeRF maps. However, it still faces limitations due to the inherent inference speed constraints of the NeRF model itself even utilizing a fast inference model like Zip-NeRF [13]. In future work, we plan to integrate and evaluate the concept of Fast Loc-NeRF with 3D renderable maps that offer faster rendering speeds, such as 3D Gaussian splitting [39].
Appendix
A Detailed Experiment Setting
For the global localization experiments in Section 4.1, we followed the initial distribution settings of Loc-NeRF [10], generating offset pose from the ground truth pose within a range of [-1m, 1m] for each axis. Particles were then uniformly sampled within a 2m $\times$ 2m $\times$ 2m cube centered on these offset poses. Additionally, we sample yaw angles from a uniform distribution within the range of [-180 , 180] and rotate the sampled particle. In comparison to our method, Loc-NeRF was configured with a batch size (ray per particle) of 32 and utilized 600 particles. We also applied $particle$ $annealing$, reducing the number of particles to 100. In the case of iNeRF [5] and PI-NeRF [6], the experimental setting follows the original methodology of each paper. The offset pose is used as the initial pose for the experiments on gradient-based optimization comparisons. For iNeRF [5], we use $batch\_size=2024$ per each gradient step, and interest region sampling as a point sampling method. Adam optimizer with an exponentially decaying learning rate ($\beta_{1}=0.9$ and $\beta_{2}=0.999$) is used to optimize the camera pose. For PI-NeRF [6], which utilized instant-NGP [14] as a map, we trained 8 scenes of LLFF dataset [36] on instant-NGP to evaluate the global localization performance. In the optimization process, 64 camera samples were utilized as parameters, and the L2 loss function was applied as the loss function. Since the update time varies depending on the rendering speed of the map, the time per update step for PI-NeRF [6] was not included in Table 1, as instant-NGP was used as the map instead of the basic NeRF [4].
B Ablation Studies
In this section, we perform a module ablation on Fast Loc-NeRF, analyzing the impact of each contribution on global localization on NeRF: coarse-to-fine framework and particle rejection weighting. Additionally, we present analyses of particle rejection weighting and the effect of the initial number of particles.
B.1 Module Ablation
@@
Fast Loc-NeRF proposes two methods: a coarse-to-fine localization framework utilizing MCL, which focuses on exploration to exploitation performing at progressively higher resolution, and a particle rejection weighting method that assigns low weights to abnormal particles outside the searching area. Table B.1 shows the results of applying each method separately. By applying a coarse-to-fine localization framework, the overall time per update step is halved while maintaining localization accuracy as shown in Table B.1. Additionally, the implementation of particle rejection weighting not only improves position accuracy by 13% but also accelerates convergence by assigning lower weights to abnormal particles, leading to a reduction in overall computation time.
Coarse-to-Fine Rejection Position Rotation Time per step (sec)
Accuracy(%) Error(cm) Accuracy(%) Error($\circ$)
52.5 0.50 65.0 5.84 0.7450
55.0 0.45 75.0 3.18 0.3836
65.0 0.33 85.0 8.12 0.5056
75.0 0.11 97.5 1.09 0.3267
Table B.1: Results of the module ablation on Fast Loc-NeRF. The table presents the position error (cm) and accuracy (%), rotation error (deg) and accuracy (%), and time per update step (sec) for different combinations of coarse-to-fine multi-scale matching and particle rejection weighting methods in LLFF dataset [36].
B.2 Analysis on Particle Rejection Weighting
Figure B.1: Uncertainty Visualization Rendered image and visualization of the uncertainty $(z_{\text{opaque}}-z_{\text{trans}})$ of each pixel from normal particles and abnormal particles on $Drjohnson$ scene of Deep Blending dataset [37]. The first row shows the properly rendered image from normal particles located on a valid region and the corresponding uncertainty for each pixel. The second row shows the strange rendered image from abnormal particles located on an invalid region and the corresponding uncertainty for each pixel, showing high values overall.
Ray Uncertainty Visualization The particle rejection weighting method is designed to filter out abnormal particles by incorporating particle rejection sampling into importance sampling. As we described in Equation 5, $z_{\text{trans}}$ and $z_{\text{opaque}}$ represent the distances along a ray at which the ray remains transparent and at which it becomes opaque, respectively. By utilizing the distance between $z_{\text{trans}}$ and $z_{\text{opaque}}$, we estimate the uncertainty of each ray. Fig B.1 visualizes the rendered images and the uncertainty of each pixel from particles. The first row shows a rendered image from normal particles on a valid region and visualizes the corresponding uncertainty for each pixel. The second row displays a rendered image and the associated uncertainty from abnormal particles on an invalid region. With normal particles, proper images well-representing scenes are rendered, while abnormal particles produce strange images. Also, the uncertainty values $(z_{\text{opaque}}-z_{\text{trans}})$ from these abnormal particles are generally high, compared to those from normal particles. This indicates that, through particle rejection weighting, abnormal particles receive lower weights, increasing the likelihood of their rejection during the resampling process. Furthermore, we show the uncertainty visualization on Fig C.1 across various scenes in the LLFF dataset [36].
Ablation Studies on Threshold of Particle Rejection Weighting
Figure B.2: Ablation results on minimum bound $\tau$. The graph shows position errors (cm) over time (sec) for different $\tau$ values evaluated on LLFF dataset [36].
@@ B.2 Analysis on Particle Rejection Weighting
Our particle rejection weighting method stabilizes by setting a minimum bound $\tau$ on the gap between the last transparent point $z_{\text{trans}}$ and the opaque point $z_{\text{opaque}}$ as $F(r(p_{j}))=\max(z_{\text{opaque}}-z_{\text{trans}},\tau)$ on Equation 6. If the gap between $z_{\text{trans}}$ and $z_{\text{opaque}}$ is very small, such as 0.01, it results in assigning too high weights to particles with rays having a small gap between these two points, rather than serving the purpose of rejecting particles on the invalid region. To prevent overweighting and ensure stability, a minimum bound $\tau$ is imposed. Fig B.2 presents the results for various minimum bounds $\tau$. As observed from Fig B.2, smaller values of $\tau$ make the particle rejection weighing more effective in reducing position errors. Additionally, when $\tau$ exceeds 0.4, the results become similar to those obtained without particle rejection weighting. The reason is that when the minimum value becomes large, the uncertainty no longer significantly influences the importance weighting process. Fig B.2 shows that our proposed particle rejection weighting effectively eliminates abnormal particles, resulting in improved global localization performance and enhanced convergence to the valid region.
B.3 Analysis on # of initial particle
Figure B.3: Ablation results on # of initial particles. The graph shows position errors (cm) over time (sec) for different # of initial particles evaluated on LLFF dataset [36].
@@ B.3 Analysis on # of initial particle
In this section, we address the number of particles used during the coarse phase of the Fast Loc-NeRF framework. Since no prior information is provided on the NeRF map $\mathcal{M}$, we initially sample the particles randomly or uniformly across the searching region. The randomly sampled particles in the initial distribution play a crucial role in the exploration process. Table B.2 illustrates the results of the global localization for varying numbers of initial particles, from 600 to 38400. Increasing the number of initial particles enhances exploration, leading to reduced position error and accuracy. However, beyond a certain number of particles, the localization inference time during the coarse phase increases exponentially. Therefore, in the coarse phase, the number of particles was set to 9600, considering the tradeoff between inference time and performance. This parameter can be adjusted based on the size of the searching space.
# of Initial Particle Position Rotation Time per step (sec)
Accuracy(%) Error(cm) Accuracy(%) Error($\circ$)
600 42.5 0.51 67.5 9.30 0.3818
1200 45.0 0.51 72.5 10.4 0.3645
2400 60.0 0.30 80.0 3.56 0.2996
4800 62.5 0.20 90.0 10.9 0.3329
9600 75.0 0.11 97.5 1.09 0.3237
19200 77.5 0.16 92.5 8.12 0.4371
38400 77.5 0.08 95.0 1.11 0.6020
Table B.2: Ablation study on the effect of the initial number of particles in Fast Loc-NeRF. The table presents the position error (cm) and rotation error (deg) for different initial particle counts.
C Implementation Detail on Zip-NeRF
Fast loc-NeRF progressively increases the resolution of rendered rays in a coarse-to-fine strategy. Our weighting method matches rendered images with observed images across various resolutions. Therefore, a 3D model that can accurately represent and render without aliasing issues at different resolutions is suitable for our proposed coarse-to-fine multiscale matching. Among various multi-scale representations [11, 12, 13, 40], we utilized Zip-NeRF [13], a grid-based approach with fast rendering speed, to serve as a map for global localization, as evaluated in Section 4.2.
To train the multi-scale representation using Zip-NeRF [13] as a map, we converted each training image into a set of four bicubically down-sampled images. By utilizing 4 sets of bicubically down-sampled images as training data, we generated a generalized model capable of effectively handling various scales. We trained the Zip-NeRF model on 8 scenes from the LLFF dataset [36] and 2 scenes from the Deep Blending dataset [37], each for 25k iterations. Zip-NeRF was implemented with two proposal MLPs and 1 final NeRF MLP. We use 64 samples for two rounds of proposal sampling each and 32 samples for the final NeRF sampling.
(a) Rendered image and uncertainty visualization from $Horns$ and $Orchichs$ scenes of LLFF dataset.
(b) Rendered image and uncertainty visualization from $Trex$ and $Room$ scenes of LLFF dataset
Figure C.1: Uncertainty Visualization Rendered image and visualization of the uncertainty $(z_{\text{opaque}}-z_{\text{trans}})$ of each pixel through heat map from normal particles and abnormal particles on LLFF dataset [36].

### v2

@@ IV-B Localization on Multi-Scale Representation
TABLE II: Comparison on LLFF dataset [36] and Deep Blender dataset [38] utilizing Zip-NeRF [13] as a map.
(a) LLFF (b) Deep Blending
Fern Leaves Flower Horns Fortress Room Orchids Trex Drjohnson Playroom Time
Position Error (cm) / Rotation Error ($\circ$)
Loc-NeRF 1.54/57.00 0.63/14.5 0.37/33.00 0.64/12.00 0.28/1.59 0.34/5.60 1.10/133.00 0.54/4.45 0.83/54.40 0.97/35.65 0.205
Ours 0.67/16.40 0.32/4.30 0.10/21.00 0.02/0.22 0.02/0.49 0.02 /0.38 1.80/84.00 0.06/0.73 0.42/34.04 0.37/33.40 0.118
Position Error $<$ 5cm / Rotation Error $<5^{\circ}$
Loc-NeRF 0.00/0.00 0.10/0.70 0.55/0.65 0.50/0.75 0.55/0.95 0.35/0.60 0.00/0.00 0.70/0.75 0.10/0.10 0.10/0.15 0.205
Ours 0.68/0.68 0.28/0.80 0.88/0.88 1.00/1.00 1.00/1.00 0.96/1.00 0.20/0.20 0.80/0.96 0.35/0.25 0.25/0.40 0.118
@@
Experiment Setup: We utilized multi-scale representation via Zip-NeRF [13] to conduct comparative analyses between Loc-NeRF and Fast Loc-NeRF. The experimental dataset consists of all 8 scenes from the LLFF dataset [36], consistent with Section IV-A, and two additional scenes from the Deep Blending dataset [38]. The Deep Blending dataset [38], unlike the LLFF dataset [36] which comprises primarily forward-facing scenes, showcases entire rooms, making it particularly suitable for demonstrating global localization for the full real world.
@@
Results We present the results of employing Zip-NeRF [13] as a map on our method and Loc-NeRF in Table II. Our approach, Fast Loc-NeRF demonstrated superior performance against baseline across most of the forward-facing scenes. While Loc-NeRF [10] had difficulty reaching a accuracy at 80% in any scene, our Fast Loc-NeRF method consistently attained this benchmark across most of the scenes. On the Deep Blending dataset, each scene representing a single room is a more challenging environment for global localization compared to the forward-facing scenes. Our method significantly outperforms Loc-NeRF shown in Table II. This enhancement suggests that our method efficiently explores extensive spatial environments.
IV-C Ablation Studies
@@
Module Ablation Fast Loc-NeRF proposes two methods: a particle rejection weighting and a coarse-to-fine localization framework. Table III shows the results of applying each method separately. By applying a coarse-to-fine localization framework, the overall time per update step is halved while maintaining localization accuracy as shown in Table III. Additionally, the implementation of particle rejection weighting not only improves position accuracy but also accelerates convergence by assigning lower weights to abnormal particles, leading to a reduction in overall computation time.
Position Rotation Time per step
Accuracy Error Accuracy Error
w/o Both 52.5 0.50 65.0 5.84 0.7450
w/o Rejection 55.0 0.45 75.0 3.18 0.3836
w/o Coarse-to-Fine 65.0 0.33 85.0 8.12 0.5056
Full 75.0 0.11 97.5 1.09 0.3267
TABLE III: Results of the module ablation on Fast Loc-NeRF.
Figure 5: Ablation studies on threshold of particle rejection weighting.
@@
Ablation Study on Particle Rejection Weighting Our particle rejection weighting method stabilizes by setting a minimum bound $\tau$ on the gap between the last transparent point $z_{\text{trans}}$ and the opaque point $z_{\text{opaque}}$ as $F(r(p_{j}))=\max(z_{\text{opaque}}-z_{\text{trans}},\tau)$ on Equation 6. Fig 5 presents the average of localization results on 8 scenes of LLFF dataset for various minimum bounds $\tau$. As observed from Fig 5, smaller values of $\tau$ make the particle rejection weighing more effective in reducing position errors. Additionally, when $\tau$ exceeds 0.4, the results become similar to those obtained without particle rejection weighting. Since the minimum value becomes $\tau$ large, the uncertainty no longer significantly influences the importance weighting process.
Figure 6: Results with various numbers of initial particles on our framework.
@@
Ablation Study on # of initial particle Since no prior information is provided on the NeRF map $\mathcal{M}$, we initially sample the particles randomly or uniformly across the searching region. Fig 6 illustrates the results of the global localization with varying numbers of initial particles. Increasing the number of initial particles improves exploration, resulting in lower position error and higher accuracy. However, beyond a certain threshold, the inference time for localization during the coarse phase increases exponentially.
V Conclusion
In summary, we introduce Fast Loc-NeRF, which enhances efficiency and accuracy in NeRF map-based global localization. To deal with particles distributed in irrelevant locations, we propose particle rejection weighting, which estimates ray uncertainty by analyzing the distribution over each ray without any additional training step. By not simply using the RGB values obtained from rendering, but also applying the uncertainty of each ray, we reduced the exploration of inappropriate space, enabling faster and more accurate localization. Furthermore, we exploit particle filtering in a coarse-to-fine manner to enable efficient localization on a NeRF map. While progressively matching particle states and measurements from low to high resolution, Fast Loc-NeRF excelerates the expansive particle weight update process. As a result, our method speeds up the localization process and improves localization accuracy by evaluating various NeRF datasets and multi-scale NeRF models.
Acknowledgements
This work was supported by Korea Evaluation Institute Of Industrial Technology (KEIT) grant funded by the Korea government(MOTIE) (No.20023455, Development of Cooperate Mapping, Environment Recognition and Autonomous Driving Technology for Multi Mobile Robots Operating in Large-scale Indoor Workspace)